# An Exact Solution Algorithm for the Bi-Level Optimization Problem of Electric Vehicles Charging Station Placement

### v1

@@ Abstract
This work addresses electric vehicle (EV) charging station placement through a bi-level optimization model, where the upper-level planner maximizes net revenue by selecting station locations under budget constraints, while EV users at the lower level choose routes and charging stations to minimize travel and charging costs. To account for range anxiety, we construct a battery-expanded network and apply a shortest path algorithm with Frank-Wolfe traffic assignment. Our primary contribution is developing the first exact solution algorithm for large scale EV charging station placement problems. We propose a Branch-and-Price-and-Cut algorithm enhanced with value function cuts and column generation. While existing research relies on heuristic methods that provide no optimality guarantees or exact algorithms that require prohibitively long runtimes, our exact algorithm delivers globally optimal solutions with mathematical certainty under a reasonable runtime. Computational experiments on the Eastern Massachusetts network (74 nodes, 248 links), the Anaheim network (416 nodes, 914 links), and the Barcelona network (110 zones, 1,020 nodes, and 2,512 links) demonstrate exceptional performance. Our algorithm terminates within minutes rather than hours, while achieving optimality gaps below 1% across all instances. This result represents a computational speedup of over two orders of magnitude compared to existing methods. The algorithm successfully handles problems with over 300,000 feasible combinations, which transform EV charging infrastructure planning from a computationally prohibitive problem into a tractable optimization task suitable for practical decision making problem for real world networks.
@@ 1 Introduction
Determining optimal locations for EV charging stations is inherently dependent on user behavior, network flow patterns, and demand distribution. Because building charging stations at every site with potential demand is infeasible due to budget constraints, planners must strategically identify locations that maximize accessibility and efficiency of usage. However, the optimal placement of charging stations is not simply a matter of available budget. It is also closely tied to user responses and the resulting traffic flow within the transportation network. Naturally, users try to minimize their travel and charging costs, meanwhile their collective choices influence network utilization and congestion. This paper addresses the charging station deployment problem using a bi-level optimization framework. At the upper level, a network planner seeks to design a profitable charging network by selecting station locations within a fixed budget and estimating revenue from user charging behavior. Moreover, EV users act as rational agents that minimize travel and charging costs, which depend on the network configuration. The interplay between these two decision making layers requires a two-level model that captures both planner objectives and user equilibrium (UE) responses.
@@
Many papers have explored the charging station placement for EVs using a bi-level framework [doi:10.1287/trsc.2021.0494, 10122631, 10034431, 9690622, 9944904]. 9944904 plan urban roads, EV charging stations, and the power grid by a bi-level model formulation where the upper level (UL) is investment decisions and the lower level (LL) is EV UE routing. They convert the entire model to a single mixed integer linear program (MILP) via KKT conditions and an improved Big-$M$ linearization, then solve it exactly with a solver. Their formulation relies on very large $M$ constants. Large $M$ MILPs can still suffer weak relaxations and long solve times on bigger networks. Even the 41 nodes/56-bus of their case study is labeled “large scale”; the authors note that adding expressways or finer zoning could induce larger computational burdens. The paper itself hints that decomposition (e.g., Benders) or scenario reduction would be needed for bigger instances, but these techniques are not yet implemented.
@@
In another study, https://doi.org/10.1155/2017/4252946 formulates a bi-level problem where the upper level planner selects some roadway link candidates to host chargers to maximize battery electric vehicle (BEV) flow, while a LL computes a stochastic UE (SUE) restricted by path distance in which BEVs are limited by driving range. Based on the paper, the resulting mixed integer nonlinear problem (MINP) is intractable for exact solvers, and the authors propose a heuristic loop. They demonstrate the method on two benchmark road networks: Nguyen-Dupuis (13 nodes, 19 links, 4 O-D pairs) and Sioux Falls (24 nodes, 76 links, 576 O-D pairs). While their algorithm is fast on small graphs, this approach does not offer an optimality guarantee. It can stall at a local maximum of covered BEV flow and provides no bound on how close that solution is to the true optimum. Similarly, [Tran10122021] used a purely heuristic called Cross-Entropy Method (CEM). UL chooses nodes to minimize the sum of the stations’ installation costs and the value of time spent traveling. At the same time, LL assigns mixed EV traffic via a multi class UE that respects EV range. Again, it is shown that the algorithm converges in their test networks, but it is not guaranteed to converge to the global optimum. The algorithm provides no optimality gap or confidence interval and works only on small to medium networks.
@@
In contrast to these earlier studies, our work provides the first exact solution methodology that scales to metropolitan scale road networks. We formulate the charging station placement problem as a MINP identical in fidelity to the small network models above, but we solve it directly without surrogate objectives, path enumeration, or heuristic sampling. The resulting algorithm delivers global optimality on networks two orders of magnitude larger than those in prior studies within practical runtimes, and it produces optimality gaps of less than 1% in all tested instances. Hence, our contribution is twofold: we close the optimality gap left by heuristic approaches and we demonstrate tractability on very large, real world networks previously considered beyond reach. Building on insights from recent literature, the UL decision involves binary location variables and station cost constraints, while the LL subproblem models traffic assignment and charging flows using a convex UE formulation. We propose an exact solution approach based on the high point relaxation (HPR) and value function cuts to globally approximate the optimal bi-level solution. This formulation enables network planners to anticipate user behavior and optimize system design in a way that balances profitability and user accessibility.
@@ 2.1 Discrete Traffic Network Design Problem
One of the first papers on DNDP problems was leblanc1975algorithm, who proposed a model with binary decision variables for link construction under budget constraints. The problem minimize network congestion through system optimal (SO) flows. They obtained lower bounds using HPR, and replaced UE with system optimum traffic assignment while optimistically assuming all unfixed candidate links are built and ignoring budget constraints. Farvaresh2013 argued that fixing unfixed variables at 1 with system optimal flows creates inefficient lower bounds with high gaps. They implemented a B&B with fathoming processes and formulated a mixed integer convex program solved exactly with outer approximation (OA) loops, which yields tighter lower bounds by keeping both budget and binary status of undecided links. They also revised the path-based formulation into a link-node multi-commodity network. So they eliminated explicit path enumeration and enabled modern assignment routines.
@@
WANG201342 solved the harder multi capacity DNDP, where each candidate link may receive 0, 1, or more additional lanes. They keep leblanc1975algorithm’s insight on SO routing gives an admissible lower bound, but embed it in two global optimization schemes that avoid the looseness Farvaresh2013 still inherits from switching to SO only after branching. Their first scheme, SO relaxation, sorts all designs by increasing SO travel time and successively evaluates them under UE. The second, tighter UE reduction, adds the Beckmann–McGuire–Winsten objective of the UE problem as a convex inequality to the SO model, thereby shrinking the relaxation without losing validity. https://doi.org/10.1111/mice.12224 replaced Farvaresh2013’s OA lower bound with an embedded Benders decomposition: at every B&B node, the SO DNDP is written as a MINP and decomposed into a convex SO traffic assignment subproblem and an MILP master. The dual produced by the subproblem yields a markedly tighter bound than the previous algorithms. These refinements allow the first exact treatment of a real scale, multimodal, multiclass network (Winnipeg: 943 nodes, 3,075 links, 20 projects).
@@
ReyLevin2024 reformulate the SO relaxation of the DNDP in a path space and exploit two structural facts that earlier exact methods left untapped: (i) the total system travel time objective is link-separable and therefore amenable to per link OA; and (ii) the exponential path set can be navigated on demand via column generation (CG), so no path enumeration is needed. These relaxations feed a single tree branch and price and cut (BPC) algorithm, where lower bounds come from the evolving LP (restricted master), and upper bounds are obtained by solving a UE whenever a node becomes feasible regarding the budget. Compared with [Farvaresh2013]’s link-based OA B&B, the path-based master removes multi commodity link constraints and scales the OA cut only where needed. Unlike https://doi.org/10.1111/mice.12224’s B&B–Benders, which still solves a nonlinear SO subproblem at every node, the BPC bound is a linear LP and thus faster per evaluation.
@@ 2.2 EV charging station problem
Many papers have considered the problem of choosing locations for the charging station for EVs without considering the response of the EV users [doi:10.3141/2252-12, 10406825, 8388731, 10703145, 9713435]. For example, doi:10.3141/2252-12 estimated the day and night charge demand of EVs through census and employment regressions and solved a mixed integer problem (MIP) capable of delivering maximum coverage to site chargers for maximum coverage within a certain radius. However, by treating drivers as fixed demand points and omitting any UE feedback, the model cannot capture route choice and EV user response to station placement. 10406825 solved EV charger siting and sizing under uncertain range and demand as a two stage stochastic MILP on 1,079 Pennsylvania demand nodes. The model treats vehicles as fixed demands with fictitious travel penalties, so it neglects route choice, congestion and other UE responses that a bi-level formulation would capture.
@@
Some papers have considered a bi-level formulation and capturing the user’s response [doi:10.1287/trsc.2021.0494, 10122631, 10034431, 9690622, 9387405, 10196503, 8064175, https://doi.org/10.1155/2017/4252946, HE2015227, 10422114, 8064175]. They can be divided into two categories, on whether they propose exact or heuristic solution algorithms. doi:10.1287/trsc.2021.0494 addressed strategic EV charging station location and sizing under stochastic flows and congestion using bi-level optimization. The UL minimized infrastructure costs while ensuring probabilistic service level constraints, while the LL represented EV users minimizing route lengths. They simplified the bi-level model into a single-level MILP with M/M/c queuing and developed an exact Benders’ decomposition solution. However, the algorithm only guarantees global optimality on small networks. For larger cases, the MILP becomes intractable and results may be far from optimal with unknown gaps. 10122631 proposed a bi-level model for EV charging station placement and direct current fast charger in highway networks, with the UL minimizing total annualized costs and the LL addressing UE traffic assignment and power distribution network operations. They converted this to a single-level MILP using KKT conditions, ensuring global optimality. Testing was limited to a modest 3-city network with 6 OD pairs and 14 pregenerated paths, leaving practical tractability on larger instances with thousands of potential paths unclear.
@@
10034431 developed a hierarchical EV charging network design model with dynamic pricing. Their problem minimize facility and operating costs while maximizing operator revenue balanced against user travel and charging expenses. They reformulated the bi-level problem into a single-level model using KKT conditions and solved it with an iterative active set heuristic alternating location fixes with Macroscopic-Fundamental-Diagram (MFD) updates for travel times. The results on a network with 12 nodes showed a gap within 4.5% of the exact solution, while on a 42 nodes campus network they achieved a 0.43% gap. Without global optimality guarantees, this solution is a near-optimal heuristic rather than an exact solver. In another study, 9690622 developed a hierarchical model for EV charging infrastructure design and operational management. The UL minimize deployment costs and maximize revenue through dynamic pricing, and the LL minimize user travel and charging expenses UE conditions. They implemented Mitsos’s global optimization bounding scheme, which theoretically reaches global optimum but practically stopped at 4.58% and 2.89% gaps on test cases. This yields a solution that is near optimal. The algorithm required tens of hours for networks under 350 links, which shows intractability for larger networks without additional strategies.
@@
HE2015227 framed public charger siting as a bi-level problem with budget constraint. The UL choose station locations and the LL use tour-based equilibrium for EV drivers which select routes and recharge stops, based on battery range and risk preferences. They reformulated this as a path-based program solved with a genetic algorithm (GA) coupled to CG for the equilibrium subproblem. Since the GA is heuristic without optimality guarantees and testing was limited to the modest Sioux Falls network (24 nodes, 76 links), scalability and solution quality on larger networks remain uncertain. 10422114 proposed a bi-level model for electric autonomous vehicles with an UL modified p-median locating fixed fast charging stations to minimize flow weighted access distance, and a LL SO traffic assignment with stochastic charging demand minimizing total travel, charging, and queuing costs. They solved this with an iterative heuristic alternating traffic re-assignment and station re-siting, which relax the nonlinear LL via generalized Lagrangian and Frank Wolfe approximation. Testing only on the small Nguyen-Dupuis case (13 nodes, 19 links) leaves scalability and global optimality unproven. The authors themselves call for validation on larger networks.
@@
Our contribution addresses critical limitations identified in prior studies by explicitly modeling EV users’ responses via a UE LL within a bi-level framework. Unlike earlier exact algorithms that lose tractability beyond small networks or heuristic methods lacking optimality guarantees, we develop an exact solution algorithm that achieves proven global optimality in a reasonable runtime. Furthermore, we demonstrate its practical scalability by solving substantially larger, realistic networks while maintaining very small optimality gaps, a feat not previously accomplished in the existing literature.
@@ 3 Formulation
We consider a bi-level formulation. The UL decision maker (network planner) designs an EV charging network and tries to maximize their revenue. At the same time, the LL represents the users, who choose charging stations to minimize their travel and charging costs, given the network design set by the planner. Because the planner’s revenue depends on the flow each station ultimately attracts, the UL objective is evaluated through the LL equilibrium. The planner’s revenue depends on the users’ choices, which will create a hierarchical optimization structure where the UL must solve the LL problem to evaluate any potential network design. For the route choice behavior of the LL problem, we assume that users follow UE principles, which means users seek the path with the minimum cost. Most papers whose model is traffic assignment make the same assumptions, too. The mathematical formulation of UE can be solved as a convex program because our model satisfies the necessary conditions: 1— We have assumed that users have perfect knowledge of the path cost; 2— The travel cost on a given link depends only on the flow on that link, and the travel cost functions are positive and monotonically increasing; 3— Each user selects the path with the least travel cost between their origin and destination. Under these assumptions, the UE condition is characterized by the principle that all used routes connecting the same origin-destination (OD) pair have equal and minimal travel cost. Unused routes may have higher travel costs, and routes connecting different OD pairs may have different travel costs. A stable equilibrium is achieved when no user can reduce travel time by unilaterally changing routes. This formulation of the UE condition, first proposed by doi:10.1680/ipeds.1952.11259, has been widely applied in transportation research. By solving the UE traffic assignment problem (TAP), we can output the route flows and each charging station flow that minimize travel cost in equilibrium across the network.

### v2

@@ Abstract
This work addresses electric vehicle (EV) charging station placement through a bi-level optimization model, where the upper-level planner maximizes net revenue by selecting station locations under budget constraints, while EV users at the lower level choose routes and charging stations to minimize travel and charging costs. To account for range anxiety, we construct a battery-expanded network and apply a shortest path algorithm with Frank-Wolfe traffic assignment. Our primary contribution is developing the first exact solution algorithm for large scale EV charging station placement problems. We propose a Branch-and-Price-and-Cut algorithm enhanced with value function cuts and column generation. While existing research relies on heuristic methods that provide no optimality guarantees or exact algorithms that require prohibitively long runtimes, our exact algorithm delivers globally optimal solutions with mathematical certainty under a reasonable runtime. Computational experiments on the Eastern Massachusetts network (74 nodes, 248 links), the Anaheim network (416 nodes, 914 links), and the Barcelona network (110 zones, 1,020 nodes, and 2,512 links) demonstrate exceptional performance. Our algorithm terminates within minutes, while achieving optimality gaps below 1% across all instances. Based on comparisons with published results with similar problems, our algorithm solves substantially larger networks with tighter optimality gaps in less computational time than existing methods. The algorithm successfully handles problems with over 300,000 feasible combinations, which transform EV charging infrastructure planning from a computationally prohibitive problem into a tractable optimization task suitable for practical decision making problem for real world networks.
@@ 1 Introduction
Determining optimal locations for EV charging stations is a critical infrastructure planning challenge that depends on multiple interdependent factors, including user behavior, network flow patterns, demand distribution, and the broader energy ecosystem. Beyond serving transportation needs, strategically located stations can also reduce the economic and safety concerns associated with fuel transportation [9]. Since building charging stations at every site with potential demand is infeasible due to budget constraints, planners must strategically identify locations that maximize accessibility and efficiency of usage. However, the optimal placement of charging stations is not simply a matter of available budget. It is also closely tied to user responses and the resulting traffic flow within the transportation network. Naturally, users try to minimize their travel and charging costs, meanwhile their collective choices influence network utilization and congestion. This paper addresses the charging station deployment problem using a bi-level optimization framework. At the upper level, a network planner seeks to design a profitable charging network by selecting station locations within a fixed budget and estimating revenue from user charging behavior. Moreover, EV users act as rational agents that minimize travel and charging costs, which depend on the network configuration. The interplay between these two decision making layers requires a two-level model that captures both planner objectives and user equilibrium (UE) responses.
@@
Many papers have explored the charging station placement or for EVs using a bi-level framework [13, 27, 20, 19, 16]. Li et al. [16] plan urban roads, EV charging stations, and the power grid by a bi-level model formulation where the upper level (UL) is investment decisions and the lower level (LL) is EV UE routing. They convert the entire model to a single mixed integer linear program (MILP) via KKT conditions and an improved Big-$M$ linearization, then solve it exactly with a solver. Their formulation relies on very large $M$ constants. Large $M$ MILPs can still suffer weak relaxations and long solve times on bigger networks. Even the 41 nodes/56-bus of their case study is labeled “large scale”; the authors note that adding expressways or finer zoning could induce larger computational burdens. The paper itself hints that decomposition (e.g., Benders) or scenario reduction would be needed for bigger instances, but these techniques are not yet implemented.
@@
In another study, Jing et al. [12] formulates a bi-level problem where the upper level planner selects some roadway link candidates to host chargers to maximize battery electric vehicle (BEV) flow, while a LL computes a stochastic UE (SUE) restricted by path distance in which BEVs are limited by driving range. Based on the paper, the resulting mixed integer nonlinear problem (MINP) is intractable for exact solvers, and the authors propose a heuristic loop. They demonstrate the method on two benchmark road networks: Nguyen-Dupuis (13 nodes, 19 links, 4 O-D pairs) and Sioux Falls (24 nodes, 76 links, 576 O-D pairs). While their algorithm is fast on small graphs, this approach does not offer an optimality guarantee. It can stall at a local maximum of covered BEV flow and provides no bound on how close that solution is to the true optimum. Similarly, [25] used a purely heuristic called Cross-Entropy Method (CEM). UL chooses nodes to minimize the sum of the stations’ installation costs and the value of time spent traveling. At the same time, LL assigns mixed EV traffic via a multi class UE that respects EV range. Again, it is shown that the algorithm converges in their test networks, but it is not guaranteed to converge to the global optimum. The algorithm provides no optimality gap or confidence interval and works only on small to medium networks.
@@
In contrast to these earlier studies, our work provides the first exact solution methodology that scales to metropolitan scale road networks. We formulate the charging station placement problem as a MINP identical in fidelity to the small network models above, but we solve it directly without surrogate objectives, path enumeration, or heuristic sampling. The resulting algorithm delivers global optimality on networks two orders of magnitude larger than those in prior studies within practical runtimes, and it produces optimality gaps of less than 1% in all tested instances. Hence, our contribution is twofold: we close the optimality gap left by heuristic approaches and we demonstrate tractability on very large, real world networks previously considered beyond reach. Building on insights from recent literature, the UL decision involves binary location variables and station cost constraints, while the LL subproblem models traffic assignment and charging flows using a convex UE formulation. We propose a Branch-and-Price-and-Cut (BPC) algorithm enhanced with value function cuts and column generation, built upon the high point relaxation (HPR), to solve the bi-level problem to global optimality. This formulation enables network planners to anticipate user behavior and optimize system design in a way that balances profitability and user accessibility.
@@ 2.1 Discrete Traffic Network Design Problem
One of the first papers on DNDP problems was LeBlanc [15], who proposed a model with binary decision variables for link construction under budget constraints. The problem minimize network congestion through system optimal (SO) flows. They obtained lower bounds using HPR, and replaced UE with system optimum traffic assignment while optimistically assuming all unfixed candidate links are built and ignoring budget constraints. Farvaresh and Sepehri [6] argued that fixing unfixed variables at 1 with system optimal flows creates inefficient lower bounds with high gaps. They implemented a B&B with fathoming processes and formulated a mixed integer convex program solved exactly with outer approximation (OA) loops, which yields tighter lower bounds by keeping both budget and binary status of undecided links. They also revised the path-based formulation into a link-node multi-commodity network. So they eliminated explicit path enumeration and enabled modern assignment routines.
@@
Wang et al. [26] solved the harder multi capacity DNDP, where each candidate link may receive 0, 1, or more additional lanes. They keep LeBlanc [15]’s insight on SO routing gives an admissible lower bound, but embed it in two global optimization schemes that avoid the looseness Farvaresh and Sepehri [6] still inherits from switching to SO only after branching. Their first scheme, SO relaxation, sorts all designs by increasing SO travel time and successively evaluates them under UE. The second, tighter UE reduction, adds the Beckmann–McGuire–Winsten objective of the UE problem as a convex inequality to the SO model, thereby shrinking the relaxation without losing validity. Bagloee et al. [1] replaced Farvaresh and Sepehri [6]’s OA lower bound with an embedded Benders decomposition: at every B&B node, the SO DNDP is written as a MINP and decomposed into a convex SO traffic assignment subproblem and an MILP master. The dual produced by the subproblem yields a markedly tighter bound than the previous algorithms. These refinements allow the first exact treatment of a real scale, multimodal, multiclass network (Winnipeg: 943 nodes, 3,075 links, 20 projects).
@@
Rey and Levin [22] reformulate the SO relaxation of the DNDP in a path space and exploit two structural facts that earlier exact methods left untapped: (i) the total system travel time objective is link-separable and therefore amenable to per link OA; and (ii) the exponential path set can be navigated on demand via column generation (CG), so no path enumeration is needed. These relaxations feed a single tree BPC algorithm, where lower bounds come from the evolving LP (restricted master), and upper bounds are obtained by solving a UE whenever a node becomes feasible regarding the budget. Compared with [6]’s link-based OA B&B, the path-based master removes multi commodity link constraints and scales the OA cut only where needed. Unlike Bagloee et al. [1]’s B&B–Benders, which still solves a nonlinear SO subproblem at every node, the BPC bound is a linear LP and thus faster per evaluation.
@@ 2.2 EV charging station problem
Many papers have considered the problem of choosing locations for the charging station for EVs without considering the response of the EV users [8, 14, 5, 23, 17, 24]. For example, Frade et al. [8] estimated the day and night charge demand of EVs through census and employment regressions and solved a mixed integer problem (MIP) capable of delivering maximum coverage to site chargers for maximum coverage within a certain radius. However, by treating drivers as fixed demand points and omitting any UE feedback, the model cannot capture route choice and EV user response to station placement. Klein and Trautmann [14] solved EV charger siting and sizing under uncertain range and demand as a two stage stochastic MILP on 1,079 Pennsylvania demand nodes. The model treats vehicles as fixed demands with fictitious travel penalties, so it neglects route choice, congestion and other UE responses that a bi-level formulation would capture.
@@
Some papers have considered a bi-level formulation and capturing the user’s response [13, 27, 20, 19, 7, 18, 29, 12, 10, 11, 29]. They can be divided into two categories, on whether they propose exact or heuristic solution algorithms. Kınay et al. [13] addressed strategic EV charging station location and sizing under stochastic flows and congestion using bi-level optimization. The UL minimized infrastructure costs while ensuring probabilistic service level constraints, while the LL represented EV users minimizing route lengths. They simplified the bi-level model into a single-level MILP with M/M/c queuing and developed an exact Benders’ decomposition solution. However, the algorithm only guarantees global optimality on small networks. For larger cases, the MILP becomes intractable and results may be far from optimal with unknown gaps. Wang et al. [27] proposed a bi-level model for EV charging station placement and direct current fast charger in highway networks, with the UL minimizing total annualized costs and the LL addressing UE traffic assignment and power distribution network operations. They converted this to a single-level MILP using KKT conditions, ensuring global optimality. Testing was limited to a modest 3-city network with 6 OD pairs and 14 pregenerated paths, leaving practical tractability on larger instances with thousands of potential paths unclear.
@@
Mirheli and Hajibabai [20] developed a hierarchical EV charging network design model with dynamic pricing. Their problem minimize facility and operating costs while maximizing operator revenue balanced against user travel and charging expenses. They reformulated the bi-level problem into a single-level model using KKT conditions and solved it with an iterative active set heuristic alternating location fixes with Macroscopic-Fundamental-Diagram (MFD) updates for travel times. The results on a network with 12 nodes showed a gap within 4.5% of the exact solution, while on a 42 nodes campus network they achieved a 0.43% gap. Without global optimality guarantees, this solution is a near-optimal heuristic rather than an exact solver. In another study, Mirheli and Hajibabai [19] developed a hierarchical model for EV charging infrastructure design and operational management. The UL minimize deployment costs and maximize revenue through dynamic pricing, and the LL minimize user travel and charging expenses UE conditions. They implemented Mitsos’s global optimization bounding scheme, which theoretically reaches global optimum but practically stopped at 4.58% and 2.89% gaps on test cases. This yields a solution that is near optimal. The algorithm required tens of hours for networks under 350 links, which shows intractability for larger networks without additional strategies.
@@
He et al. [10] framed public charger siting as a bi-level problem with budget constraint. The UL choose station locations and the LL use tour-based equilibrium for EV drivers which select routes and recharge stops, based on battery range and risk preferences. They reformulated this as a path-based program solved with a genetic algorithm (GA) coupled to CG for the equilibrium subproblem. Since the GA is heuristic without optimality guarantees and testing was limited to the modest Sioux Falls network (24 nodes, 76 links), scalability and solution quality on larger networks remain uncertain. Hu et al. [11] proposed a bi-level model for electric autonomous vehicles with an UL modified p-median locating fixed fast charging stations to minimize flow weighted access distance, and a LL SO traffic assignment with stochastic charging demand minimizing total travel, charging, and queuing costs. They solved this with an iterative heuristic alternating traffic re-assignment and station re-siting, which relax the nonlinear LL via generalized Lagrangian and Frank Wolfe approximation. Testing only on the small Nguyen-Dupuis case (13 nodes, 19 links) leaves scalability and global optimality unproven. The authors themselves call for validation on larger networks.
@@
Our contribution addresses critical limitations identified in prior studies by explicitly modeling EV users’ responses via a bi-level framework which uses UE as the LL problem. Unlike earlier exact algorithms that lose tractability beyond small networks or heuristic methods lacking optimality guarantees, we develop an exact solution algorithm that achieves proven global optimality in a reasonable runtime. Furthermore, we demonstrate its practical scalability by solving substantially larger, realistic networks while maintaining very small optimality gaps, a feat not previously accomplished in the existing literature.
@@ 3 Formulation
We consider a bi-level formulation. The UL decision maker (network planner) designs an EV charging network and tries to maximize their revenue. At the same time, the LL represents the users, who choose charging stations to minimize their travel and charging costs, given the network design set by the planner. Because the planner’s revenue depends on the flow each station ultimately attracts, the UL objective is evaluated through the LL equilibrium. The planner’s revenue depends on the users’ choices, which will create a hierarchical optimization structure where the UL must solve the LL problem to evaluate any potential network design. For the route choice behavior of the LL problem, we assume that users follow UE principles, which means users seek the path with the minimum cost. Most papers whose model is traffic assignment make the same assumptions, too. The mathematical formulation of UE can be solved as a convex program because our model satisfies the necessary conditions: 1— We have assumed that users have perfect knowledge of the path cost; 2— The travel cost on a given link depends only on the flow on that link, and the travel cost functions are positive and monotonically increasing; 3— Each user selects the path with the least travel cost between their origin and destination. Under these assumptions, the UE condition is characterized by the principle that all used routes connecting the same origin-destination (OD) pair have equal and minimal travel cost. Unused routes may have higher travel costs, and routes connecting different OD pairs may have different travel costs. A stable equilibrium is achieved when no user can reduce travel time by unilaterally changing routes. This formulation of the UE condition, first proposed by Wardrop [28], has been widely applied in transportation research. By solving the UE traffic assignment problem (TAP), we can output the route flows and each charging station flow that minimize travel cost in equilibrium across the network.